OpenStudy (anonymous):

PLEASE HELP I REALLY NEED TO UNDERSTAND THIS!!!! Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = - x - 6 y = 3x - 14 y = - x - 14 y = 3x - 6

OpenStudy (anonymous):

Please help!!! && an explanation would be great!!

OpenStudy (anonymous):

@shifuyanli

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

first we change it into the original form y=mx+b

Mehek (mehek14):

so it becomes -y = -3x -3 since y is negative we turn it into positive so the sign changes for all of them y = 3x + 3

Mehek (mehek14):

the slope has to be the same so look for answer with 3x but the y-int can be any number

Mehek (mehek14):

we have 2 that could be the answer b or d

OpenStudy (anonymous):

Oh okay, that makes sense...

Mehek (mehek14):

so now we sub in the points 4, -2

OpenStudy (anonymous):

I was thinking b but not sure

Mehek (mehek14):

-2 = 3*4 - 14 3*4 = 12 -2 = 12 - 14 -2 = -2 it is B

Mehek (mehek14):

remember parallel lines have same slope but differ y-int

Mehek (mehek14):

so B is the answer

OpenStudy (anonymous):

OHH okay that makes sense!!

OpenStudy (anonymous):

Thanks so much!!!

Mehek (mehek14):

yw if you have 2 answer choices that are alike, just sub the points

OpenStudy (anonymous):

Okay, thank you! I will keep this in mind, :)

OpenStudy (anonymous):

One other question? @Mehek14 :/ How would I do it if instead of parallel it was perpendicular?

Mehek (mehek14):

the slope would be the reciprocal since the parallel slope is 3x for perpendicular it would be \(\dfrac{1}{3}x\)

OpenStudy (anonymous):

Thank you so much for all ya help!!